OpenStudy (anonymous):

prove the following trigonometric identities. Include reason with calculation: (secx+tanx)(1-sinx/cosx)=1

myininaya (myininaya):

lol oops i read the question wrong

myininaya (myininaya):

\[(\frac{1}{cosx}+\frac{sinx}{cosx})(1-\frac{sinx}{cosx})=\frac{1}{cosx}-\frac{1}{cosx}\frac{sinx}{cosx}+\frac{sinx}{cosx}-\frac{sinx}{cosx}\frac{sinx}{cosx}\] \[\frac{1}{cosx}-\frac{sinx}{\cos^2x}+\frac{sinx}{cosx}-\frac{\sin^2x}{\cos^2x}\] lets find a common denominator \[\frac{cosx-sinx+sinxcosx-\sin^2x}{\cos^2x}\] \[\frac{-(\sin^2x-sinxcosx)+cosx-sinx}{1-\sin^2x}=\frac{-sinx(sinx-cosx)+cosx-sinx}{(1-sinx)(1+sinx)}\] \[=\frac{sinx(cosx-sinx)+(cosx-sinx)}{(1-sinx)(1+sinx)}=\frac{(cosx-sinx)(sinx+1)}{(1-sinx)(1+sinx)}\] \[\frac{cosx-sinx}{1-sinx}\]

myininaya (myininaya):

i'm not sure if this does =1

myininaya (myininaya):

nope it does not

myininaya (myininaya):

i'm going to sleep now maybe you meant to say (secx+tanx)((1-sinx)/cosx)=1